# With Ears to See and Eyes to Hear: Sound Symbolism Experiments with Multimodal Large Language Models

### v1

@@ Abstract
Recently, Large Language Models (LLMs) and Vision Language Models (VLMs) have demonstrated aptitude as potential substitutes for human participants in experiments testing psycholinguistic phenomena. However, an understudied question is to what extent models that only have access to vision and text modalities are able to implicitly understand sound-based phenomena via abstract reasoning from orthography and imagery alone. To investigate this, we analyse the ability of VLMs and LLMs to demonstrate sound symbolism (i.e., to recognise a non-arbitrary link between sounds and concepts) as well as their ability to “hear” via the interplay of the language and vision modules of open and closed-source multimodal models. We perform multiple experiments, including replicating the classic Kiki-Bouba and Mil-Mal shape and magnitude symbolism tasks, and comparing human judgements of linguistic iconicity with that of LLMs. Our results show that VLMs demonstrate varying levels of agreement with human labels, and more task information may be required for VLMs versus their human counterparts for in silico experimentation. We additionally see through higher maximum agreement levels that Magnitude Symbolism is an easier pattern for VLMs to identify than Shape Symbolism, and that an understanding of linguistic iconicity is highly dependent on model size.
@@ 1 Introduction
Sound symbolism refers to a perceived similarity between speech sounds and the conceptual meanings of the words they comprise. Evidence of this can be found in linguistic devices such as onomatopoeia (e.g., “bang”, “shriek”, and “bellow”), where a word imitates the concept it describes via its phonetic form. The ability of Large Language Models (LLMs) and Vision Language Models (VLMs) to reflect a sense of sound symbolism would therefore suggest that these models are capable of acquiring “phonetic” knowledge indirectly through only the written orthographic form of a language via patterns of grapheme combinations and meta-level textual discussion of sound in training data, which has implications for the potential future use of LLM/VLMs in perceptual studies usually reserved for humans (Jain et al., 2023; Dillion et al., 2023; Aher et al., 2023). In this work, we further explore the capability of LLMs/VLMs to demonstrate human-like characteristics in a range of psycholinguistic perceptual tests investigating 3 main areas of sound symbolism: (1) Shape Symbolism (i.e., the Kiki-Bouba effect, Ramachandran and Hubbard, 2001), where a forced choice must be made between two pseudowords as to which is the most appropriate to describe shapes and entities that are spiky or rounded; (2) Magnitude Symbolism (i.e., the Mil-Mal Effect, Sapir, 1929), a similar test to (1), but where the entities are small or large (rather than spiky or rounded); and (3) Iconicity Rating Winter et al. (2023), where LLMs are asked to rate a series of English words on their perceived “iconicity” (i.e., to what extent a word’s form is perceived to be analogous to the concept or entity it describes).
@@
By extending these experiments to LLMs/VLMs, our study aims to shed light on the processes underlying multimodal perception in language models. Moreover, the presence of sound symbolism in such models could inform the development of more effective natural language processing algorithms, aiding tasks such as sentiment analysis, emotion recognition, and content generation that take into account more abstract layers of human reasoning and perception such as abstract connotations between words rather than semantics alone Manzoor et al. (2023). An understanding of sound symbolism also has the potential to have a profound effect on creative generation, including language forms such as poetry and narratives (and their accompanying illustrations). Additionally, sound symbolism is a prevalent strategy used in marketing products to create desirable associations in potential customers, and LLMs capable of understanding this phenomenon could be used as pilot testing before using focus groups to reduce time and monetary costs (Ketron and Spears, 2021; Motoki et al., 2020; Spence, 2012).
We summarise our main contributions as follows: [^1] Our paper title is inspired by: https://en.wikipedia.org/wiki/With_Ears_to_See_and_Eyes_to_Hear [^2] We will release all of our code on GitHub.
@@ 2 Related Works
Recently, Alper and Averbuch-Elor (2023) have investigated the ability of language models to exhibit the Kiki/Bouba effect using CLIP and Stable Diffusion by generating images from sound-symbolic prompts, and provide positive evidence for this association to be present. We build upon this work in §3 by introducing a wider range of VLMs in a forced naming task for “real” entities as opposed to abstract shapes, and additionally extend this to magnitude symbolism via the Mil-Mal test in §4. We further differentiate our work by focussing on the task of assigning pseudowords to provided visual stimuli, in contrast to Alper and Averbuch-Elor (2023) who investigate the effects of different pseudowords on the outputs of image generation models.
Additionally, in recent times large-scale efforts have also been made to collect ratings on linguistic iconicity (i.e., the level of symbolism a particular word has), with Winter et al. (2023) collecting ratings of over 14k English words. Some effort has been made to analyse whether or not similar ratings would be assigned by an LLM, where Trott (2024) used GPT-4 and reports a moderate positive correlation across ratings. We build upon this work in §5 by introducing a wider range of VLMs, including open-source alternatives. Numerous computational works in NLP have investigated other aspects of iconicity and sound symbolism, with Abramova and Fernández (2016) investigating the word embeddings of different aspects of morphology in relation to symbolism (see also Yamshchikov et al., 2019; Liu et al., 2018). Additionally, Sabbatino et al. (2022) investigated the emotional intensity of nonsense words using NLP methods to determine which phoneme combinations were most responsible. Several works in similar areas have also exemplified the ability of LLMs to demonstrate perceptual behaviour akin to humans and the potential for these models to replace human participants in pilot studies, as well as facilitating the scaling of evaluation in-silico (Jain et al., 2023; Aher et al., 2023; Dillion et al., 2023; Ramezani and Xu, 2023; Coda-Forno et al., 2023).
@@ 3.1 Methodology
We prompt DALL-E 3 (Betker et al., 2023) to generate a series of images pertaining to entities that are either “spiky” or “rounded”. Example generations can be seen in Figure 1. We use DALL-E 3 to generate examples rather than taking existing images such as the traditional representation of Kiki-Bouba in order to reduce the effects of memorisation and exhibit finer control of the physical characteristics of the presented entities. Furthermore, this allows us to further investigate the extent to which human perception of this phenomenon extends from geometric shapes to entities in the “real” world, therefore increasing ecological validity and more closely representing how LLM/VLMs may be tasked with demonstrating sound symbolism in the real world (e.g., when naming new products for marketing or new characters in a narrative). We use the prompt ”Generate an image from the following description: [spiky/rounded + noun]”, where we generate 25 spiky examples, and 25 round examples (each noun is used once per shape condition). The full list of entities can be seen in Appendix A.1. We present each prompt twice, placing each pseudoword in first or second position and then average the results, to mitigate positional biases in selections.
@@
Similar to our need to generate novel imagery to better ensure the non-memorisation of our chosen VLMs, we must avoid bias from the eponymous pseudowords (i.e., “Kiki” or “Takete” for spiked concepts, or “Bouba” and “Maluma” for round concepts). As a result of this, we consult existing sound symbolism research and select the following pseudowords that are legal in English phonotactics and imitate the same phonetic relationship that the original terms were meant to elicit. Borrowing from Occelli et al. (2013) we take: Kalika–Mabobe, Zaki-Umbu, and Tiki-Giba. From Alper and Averbuch-Elor (2023) we additionally take Kitaki-Gugagu, Hatiha-Bodubo and Penape-Gunogu. Finally, we use the original Kiki-Bouba names as a point of reference for a best-case scenario where the link is explicitly learnt from mentions within the training data.
@@
We imitate the standard human setup for the Kiki/Bouba experiment Ramachandran and Hubbard (2001) and present our VLMs with the following zero-shot prompt – ”Look at the [ENTITY] in the provided image. Out of the following two options, which name would you most likely assign to the [ENTITY]: ”[KIKI-WORD]” or ”[BOUBA-WORD]”. Respond with only your decision”. One candidate from either name category (i.e., Kiki or Bouba) is presented as outlined previously, and [ENTITY] refers to a noun used to describe the entity we wish to be named in order to direct the LLM’s attention to the correct element (i.e., the noun from the DALL-E prompt). For this experiment, we set $max\_tokens$ to 10 for the VLM responses and leave all other hyperparameters at default. We additionally provide an extended prompt we call informed, which prepends ”This task is related to the phenomenon of Sound Symbolism, which is a non-arbitrary relationship between the sound of a word and associations with its physical attributes” to give the VLMs additional task knowledge. This secondary prompt scenario is used to investigate whether human-like preferences can be encouraged from the model with additional awareness of which elements of the image to focus on.
@@ 3.2 Results
Overall, in Table 1 we see mixed results as to which model performs best, with GPT-4 showing the highest levels of agreement for the original Kiki-Bouba and the added Kalika-Mabobe, Gemini performing the best for Zaki-Umbu and Hatiha-Bodubo, and LLaVA performing the best for the remaining conditions. However, across all models, we see a general trend of low agreement with human ratings, with only a few condition/model combinations resulting in agreement above chance (50%). Regarding the introduction of the ”informed” prompt (containing additional task information), we see a general increase in agreement over the Standard condition or no change in results, indicating that once the VLMs are aware of the characteristic of interest (i.e., the shape of the entity), the VLMs are more likely to agree with human perception. However, we do see a few cases where performance decreases in the informed condition, but these are usually only minor decreases except for cases where there is already very low agreement with humans. We see varying performance across open and closed-source models. Whilst LLaVA outperforms GPT-4 and Gemini in many conditions, this is often close to chance-level agreement and may be the result of label bias, versus the closed models’ systematic disagreement. [^5] We see LLaVA variants demonstrate a clear preference for whatever pseudoword is presented in the first position. Finally, regarding our open-source LLaVA models at different sizes, we interestingly see that the 13B model outperforms the 7B and 34B models at times, most noticeably in Tiki-Giba. However, the largest 34b model performs best overall, in line with expectation (though by a small margin). Additionally, the Penape-Gunogu pair presents difficulty for our tested models, with systematic disagreement with humans by Gemini and GPT-4, suggesting the influence of additional information contained within the models’ training data as to the connotations of these pseudowords.

### v2

@@ Abstract
Recently, Large Language Models (LLMs) and Vision Language Models (VLMs) have demonstrated aptitude as potential substitutes for human participants in experiments testing psycholinguistic phenomena. However, an understudied question is to what extent models that only have access to vision and text modalities are able to implicitly understand sound-based phenomena via abstract reasoning from orthography and imagery alone. To investigate this, we analyse the ability of VLMs and LLMs to demonstrate sound symbolism (i.e., to recognise a non-arbitrary link between sounds and concepts) as well as their ability to “hear” via the interplay of the language and vision modules of open and closed-source multimodal models. We perform multiple experiments, including replicating the classic Kiki-Bouba and Mil-Mal shape and magnitude symbolism tasks and comparing human judgements of linguistic iconicity with that of LLMs. Our results show that VLMs demonstrate varying levels of agreement with human labels, and more task information may be required for VLMs versus their human counterparts for in silico experimentation. We additionally see through higher maximum agreement levels that Magnitude Symbolism is an easier pattern for VLMs to identify than Shape Symbolism, and that an understanding of linguistic iconicity is highly dependent on model size.
@@ 1 Introduction
Figure 1: Illustration of the 3 main experiments we perform. Firstly, Shape Symbolism is a binary choice between two pseudowords to best describe an object that is spiky or rounded. Magnitude Symbolism involves a binary choice between two pseudowords to best describe an object that is small or large. Finally, Iconicity involves rating the perceived iconicity of words, or how much their written/phonetic form is representative of what they describe.
@@
Sound symbolism refers to a perceived similarity between speech sounds and the conceptual meanings of the words they comprise. Evidence of this can be found in linguistic devices such as onomatopoeia (e.g., “bang”, “shriek”, and “bellow”), where a word imitates the concept it describes via its phonetic form. The ability of Large Language Models (LLMs) and Vision Language Models (VLMs) to reflect a sense of sound symbolism would therefore suggest that these models are capable of acquiring “phonetic” knowledge indirectly through only the written orthographic form of a language via patterns of grapheme combinations Loakman et al. (2024, 2023) and meta-level textual discussion of sound in training data, which has implications for the potential future use of LLM/VLMs in perceptual studies usually reserved for humans (Jain et al., 2023; Dillion et al., 2023; Aher et al., 2023). In this work, we further explore the capability of LLMs/VLMs to demonstrate human-like characteristics in a range of psycholinguistic perceptual tests investigating 3 main areas of sound symbolism: (1) Shape Symbolism (i.e., the Kiki-Bouba effect, Ramachandran and Hubbard, 2001), where a forced choice must be made between two pseudowords as to which is the most appropriate to describe shapes and entities that are spiky or rounded; (2) Magnitude Symbolism (i.e., the Mil-Mal Effect, Sapir, 1929), a similar test to (1), but where the entities are small or large (rather than spiky or rounded); and (3) Iconicity Rating Winter et al. (2023), where LLMs are asked to rate a series of English words on their perceived “iconicity” (i.e., to what extent a word’s form is perceived to be analogous to the concept or entity it describes). We illustrate these experiments in Figure 1.
@@
By extending these experiments to LLMs/VLMs, our study aims to shed light on the processes underlying multimodal perception in language models. [^1] Our paper title is inspired by the Sleeping With Sirens album of the same name: https://en.wikipedia.org/wiki/With_Ears_to_See_and_Eyes_to_Hear. Moreover, the presence of sound symbolism in such models could inform the development of more effective natural language processing algorithms, aiding tasks such as sentiment analysis, emotion recognition, and content generation that take into account more abstract layers of human reasoning and perception such as abstract connotations between words rather than semantics alone Manzoor et al. (2023). An understanding of sound symbolism also has the potential to have a profound effect on creative generation, including language forms such as poetry and narratives (and their accompanying illustrations). Additionally, sound symbolism is a prevalent strategy used in marketing products to create desirable associations in potential customers, and LLMs capable of understanding this phenomenon could be used as pilot testing before using focus groups to reduce time and monetary costs (Ketron and Spears, 2021; Motoki et al., 2020; Spence, 2012).
We summarise our main contributions as follows: [^2] We release our code and resources on GitHub: https://github.com/tylerL404/WETSAETH/.
@@ 2 Related Works
Recently, Alper and Averbuch-Elor (2023) have investigated the ability of language models to exhibit the Kiki/Bouba effect using CLIP and Stable Diffusion by generating images from sound-symbolic prompts and provide positive evidence for this association to be present. We build upon this work in §3 by introducing a wider range of VLMs in a forced naming task for “real” entities as opposed to abstract shapes, and additionally extend this to magnitude symbolism via the Mil-Mal test in §4. We further differentiate our work by focussing on the task of assigning pseudowords to provided visual stimuli, in contrast to Alper and Averbuch-Elor (2023) who investigate the effects of different pseudowords on the outputs of image generation models.
Additionally, in recent times, large-scale efforts have been made to collect ratings of linguistic iconicity (i.e., the level of symbolism a particular word has), with Winter et al. (2023) collecting ratings of over 14k English words. Some effort has been made to analyse whether or not similar ratings would be assigned by an LLM, where Trott (2024) used GPT-4 and reports a moderate positive correlation across ratings. We build upon this work in §5 by introducing a wider range of VLMs, including open-source alternatives. Numerous computational works in NLP have investigated other aspects of iconicity and sound symbolism, with Abramova and Fernández (2016) investigating the word embeddings of different aspects of morphology in relation to symbolism (see also Yamshchikov et al., 2019; Liu et al., 2018). Additionally, Sabbatino et al. (2022) investigated the emotional intensity of nonsense words using NLP methods to determine which phoneme combinations were most responsible. Several works in similar areas have also exemplified the ability of LLMs to demonstrate perceptual behaviour akin to humans and the potential for these models to replace human participants in pilot studies, as well as facilitating the scaling of evaluation in-silico (Jain et al., 2023; Aher et al., 2023; Dillion et al., 2023; Ramezani and Xu, 2023; Coda-Forno et al., 2023).
@@ 3.1 Methodology
We prompt DALL-E 3 (Betker et al., 2023) to generate a series of images pertaining to entities that are either “spiky” or “rounded”. Example generations can be seen in Figure 2. We use DALL-E 3 to generate examples rather than taking existing images such as the traditional representation of Kiki-Bouba in order to reduce the effects of memorisation and exhibit finer control of the physical characteristics of the presented entities. Furthermore, this allows us to further investigate the extent to which human perception of this phenomenon extends from geometric shapes to entities in the “real” world, therefore increasing ecological validity and more closely representing how LLM/VLMs may be tasked with demonstrating sound symbolism in the real world (e.g., when naming new products for marketing or new characters in a narrative). We use the prompt ”Generate an image from the following description: [spiky/rounded + noun]”, where we generate 25 spiky examples, and 25 round examples (each noun is used once per shape condition). The full list of entities can be seen in Appendix A.1.
@@
Similar to our need to generate novel imagery to better ensure the non-memorisation of our chosen VLMs, we must avoid bias from the eponymous pseudowords (i.e., “Kiki” or “Takete” for spiked concepts, or “Bouba” and “Maluma” for round concepts). As a result of this, we consult existing sound symbolism research papers and select the following pseudowords that are legal in English phonotactics and imitate the same phonetic relationship that the original terms were meant to elicit. Borrowing from Occelli et al. (2013) we take: Kalika–Mabobe, Zaki-Umbu, and Tiki-Giba. From Alper and Averbuch-Elor (2023) we additionally take Kitaki-Gugagu, Hatiha-Bodubo and Penape-Gunogu. Finally, we use the original Kiki-Bouba names as a point of reference for a best-case scenario where the link is explicitly learnt from mentions within the training data.
@@
We imitate the standard human setup for the Kiki/Bouba experiment Ramachandran and Hubbard (2001) and present our VLMs with the following zero-shot prompt – ”Look at the [ENTITY] in the provided image. Out of the following two options, which name would you most likely assign to the [ENTITY]: ”[KIKI-WORD]” or ”[BOUBA-WORD]”. Respond with only your decision”. One candidate from either name category (i.e., Kiki or Bouba) is presented as outlined previously, and [ENTITY] refers to a noun used to describe the entity we wish to be named in order to direct the LLM’s attention to the correct element (i.e., the noun from the DALL-E prompt). For this experiment, we set $max\_tokens$ to 10 for the VLM responses and leave all other hyperparameters at default. We additionally provide an extended prompt we call informed, which prepends ”This task is related to the phenomenon of Sound Symbolism, which is a non-arbitrary relationship between the sound of a word and associations with its physical attributes” to give the VLMs additional task knowledge. This secondary prompt scenario is used to investigate whether human-like preferences can be encouraged from the model with additional awareness of which elements of the image to focus on. We present each prompt twice, placing each pseudoword in the first or second position and then averaging the results, to mitigate positional biases in selections.
@@ 3.2 Results
Figure 3: Results of the Shape Symbolism experiments per pseudoword pair. Fleiss’ $\kappa$ Fleiss (1971) for inter-annotator agreement between humans is presented next to each pseudoword pair. Arrows indicate the direction of agreement change from the standard prompt to the informed prompt. The dashed line represents 50%, akin to chance-level agreement. Full results in table form are presented in Table 2 within Appendix A.2. In all cases, we are comparing with the human majority vote.
@@
Overall, in Figure 3 we see mixed results as to which model performs best, with GPT-4 showing the highest levels of agreement for the original Kiki-Bouba and the added Kalika-Mabobe, Gemini performing the best for Zaki-Umbu and Hatiha-Bodubo, and LLaVA performing the best for the remaining conditions. However, across all models, we see a general trend of low agreement with human ratings, with only a few condition/model combinations resulting in agreement above chance (50%). Regarding the introduction of the ”informed” prompt (containing additional task information), we see a general increase in agreement over the Standard condition or no change in results, indicating that once the VLMs are aware of the characteristic of interest (i.e., the shape of the entity), the VLMs are more likely to agree with human perception. However, we do see a few cases (e.g., Zaki/Umbu with LLaVA 7/13b, Kitaki/Gugagu with GPT-4, and Penape/Gugagu with Gemini, GPT-4, and LLaVA 13b) where performance decreases in the informed condition, but these are usually only minor decreases except for cases where there is already very low agreement with humans. We see varying performance across open and closed-source models. Whilst LLaVA outperforms GPT-4 and Gemini in many conditions, this is often close to chance-level agreement and may be the result of label bias, versus the closed models’ systematic disagreement. [^5] We see LLaVA variants demonstrate a clear preference for whatever pseudoword is presented in the first position. Finally, regarding our open-source LLaVA models at different sizes, we interestingly see that the 13B model outperforms the 7B and 34B models at times, most noticeably in Tiki-Giba. However, the largest 34b model performs best overall, in line with expectation (though by a small margin). Additionally, the Penape-Gunogu pair presents difficulty for our tested models, with systematic disagreement with humans by Gemini and GPT-4, suggesting the influence of additional information contained within the models’ training data as to the connotations of these pseudowords.
4 Magnitude Symbolism
Figure 4: Examples of “Mil”-style (tiny) and “Mal”-style (huge) generations with DALL-E 3. In total, 50 images were generated, with 25 per condition (the entities remaining constant). The ground truth is taken as the majority human vote.
4.1 Methodology
Image Dataset
Pseudowords
[…]
Task Setting
Models & Evaluation
We use the same models and evaluation protocols as in §3.1.
4.2 Results
Figure 5: Results of the Magnitude Symbolism experiments per pseudoword pair. Fleiss’ $\kappa$ Fleiss (1971) for inter-annotator agreement between humans is presented next to each pseudoword pair. Arrows indicate the direction of agreement change from the standard prompt to the informed prompt. The dashed line represents 50%, akin to chance-level agreement. Full results in table form are presented in Table 3 within Appendix A.2. In all cases, we are comparing with the human majority vote.
Overall, similar to §3 we see mixed results in Figure 5 as to which model performs best, but GPT-4 demonstrates higher levels of agreement across most conditions in a clearer pattern than what was seen in the Shape Symbolism experiment. In several cases, we see agreement that is significantly in line with human perception, with 90+agreement (e.g., GPT-4 in the Zil-Zal and Weel-Waal conditions). When comparing the standard and informed prompts, we see a much more substantial increase in performance, indicating that the VLMs fundamentally understand the relationship between sound and perceived size, but were focused on other aspects of the provided imagery when not explicitly directed towards size in the standard condition. Regarding the LLaVA models, we see the mid-sized 13B variant outperforming the 7B and 34B models in most conditions (rather than performance increasing alongside parameter count).
5 Iconicity Ratings
In this section we investigate whether LLMs demonstrate human-like associations between word forms and the entities/concepts they symbolise. [^6] We treat this as VLMs “imitating” an understanding of sound symbolism, as they of course cannot actually hear. Winter et al. (2023) present a dataset of 14k+ human judgements of iconicity to which we compare our LLM ratings (where, on a 7-point scale, “how” scored 1.3, whilst “woof” scored 6.8 due to being onomatopoeic).
5.1 Methodology
Models
Existing work by Trott (2024) has investigated whether or not GPT-4 is able to reflect human judgements and reported a Spearman correlation of 0.59. However, in trying to verify these findings, we see that our ratings differ quite strongly. For this reason, we re-run this experiment on the first 7k entries of Winter et al. (2023) and compare GPT-4 ratings as of late November 2023 with the human judgements in Winter et al. (2023) as well as the GPT-4 judgements from Trott (2024) that use an earlier version of GPT-4.777We test only on the first 7k entries due to cost. Whilst GPT-4-Turbo is markedly cheaper, this would not be a true reproduction. Please note that GPT-4o was not released at the time. In the other cases, we use the entire 14,772 entry dataset, except for GPT-3.5-Turbo, where we remove 209 words that triggered OpenAI’s safeguarding filters.
Prompting
We adopt the same prompting strategy as Trott (2024). In doing so, we use a modified version of what was presented to human participants in Winter et al. (2023). In summary, we request ratings of iconicity on a 1-7 scale, where 1 is not at all iconic, and 7 is highly iconic. The full prompt presented to the models is available in Appendix C.2.
5.2 Results
Correlations between each LLM and human judgements are presented in Table 1. We observe that the ability of LLMs to rate the iconicity of English words appears to be dependent on model size. For instance, we see no true correlation with any of our FLAN T5 models (base = 250M, XL = 3B) or our smallest LLaMA-2 variant (7B), but we begin to see significant correlations with another 7B parameter model, that being instruction-tuned Mistral-7B. When investigating our larger models, we see the 13B LLaMA-2 variant demonstrate Spearman/Pearson correlations of .379 and .381, respectively. Interestingly, however, our largest LLaMA-2 model with 70B parameters performs worse at this task than the 13B variant as seen previously, with correlations of .304/.332. Regarding OpenAI models, GPT-3.5-Turbo demonstrates a moderate correlation with humans at .420/.439. Finally, GPT-4 presents the strongest correlations. However, we observe that (on the first 7k entries), the version of GPT-4 we use (late November 2023) performs worse than the earlier version presented by Trott (2024), further demonstrating how continuously updated models also require continuous evaluation due to receiving additional training data (Spearman correlations of 0.537 vs 0.575, respectively).
Spearman Pearson
Model Corr. $p$ Corr. $p$
FLAN-T5 Base -.035 <.001 -.037 <.001
FLAN-T5 XL .000 .991 -.002 .824
Mistral-7B-Instruct .382 <.001 .377 <.001
LLaMA-2 7B -.003 .687 .005 .544
LLaMA-2 13B .379 <.001 .381 <.001
LLaMA-2 70B .304 <.001 .332 <.001
GPT-3.5-Turbo .420 <.001 .439 <.001
GPT-4 (Trott, 2024) .575 <.001 .615 <.001
GPT-4 (Ours, Nov. ’23) .537 <.001 .594 <.001
Table 1: Correlations between human ratings from Winter et al. (2023) and LLMs. GPT-4 ratings were only collected for the first 7k examples due to cost. Trott (2024) report a Spearman correlation of .590 across the entire dataset. p refers to p-value.
6 Discussion
The Source of Sound Symbolism
Firstly, due to human languages exhibiting mostly regular orthography, auditory information in speech is moderately reflected in the spellings of words via grapheme sequences (a characteristic that grapheme-to-phoneme conversion models have long exploited). Through this, text-based models are able to learn associations between grapheme sequences and semantics, based on more abstract characteristics than morpheme combinations alone, such as phonaesthemes Kaushal and Mahowald (2022). Whilst such models have no embodied understanding of sound, such statistical patterns pose a viable signal for the implicit learning of sound-based phenomena.
[…]
As a result, the relatively weak performance of our tested models could also be explained by the relative lack of sound-symbolism-heavy language in the models’ training data which is overshadowed by more prosaic language forms that do not exploit these phenomena as readily. This in turn would explain why the closed-source models we tested (e.g. GPT-4/Gemini) outperform open-source models due to the significant (assumed) differences in parameter size, allowing the larger closed models to retain more information regarding sound symbolism within the weights, in addition to being continuously updated with RLHF.
The results of our multimodal experiments additionally demonstrate that under certain conditions, VLMs show systematic disagreement with human labels, indicating the potential interference of additional knowledge contained within language model training data that influences the associations made between pseudowords and images that are not present in humans. However, it is important to note that in our experiments we compare language model selections against the majority vote or mean scores assigned by humans. Consequently, this results in a comparison to an ”ideal” human by necessity, overlooking individual differences in perception across humans (where for a decision to be ”human-like”, it has to match a choice made by any human, rather than the majority). Consequently, higher agreement levels can be observed when compared to the choices of individual humans, as inter-human agreement is not perfectly aligned in these tests.
Future Directions
7 Conclusion
Limitations
Ethics Statement
Acknowledgments
Tyler Loakman is supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1].
Visual elements in Figure 1 were taken from Flaticon. Specifically, banners (SANB), brain (Freepik), GPU (Taufik Ramadhan), ”versus” (Afif Fudin), descending/ascending (Mie Nakae), and thought bubble (Aranagraphics).
Appendix A Appendix
A.1 Implementation Details
Kiki-Bouba & Mil-Mal
For the Kiki-Bouba and Mil-Mal sound symbolism experiments, we access gemini-pro-vision via the Google Gemini API. For GPT-4, we use gpt-4-vision-preview via the OpenAI Chat Completions API. For our open-source LLaVA models at various sizes, we specifically use llava-v1.6-vicuna-7b-hf, LLaVA-v1.6-vicuna-13b-hf and LLaVA-v1.6-vicuna-34b-hf from publicly available checkpoints on Hugging Face. We use default hyperparameters for all models to test their “off-the-shelf” capability. Human participants were shown generated imagery with the binary pseudo-word options via Google Forms. The order of image presentation was randomised, whilst the order of the pseudowords was kept static. Separate Google Forms were used for each pseudoword pair. Participants were able to complete the forms at their own pace within a period of approximately 2 weeks.
Iconicity Ratings
Entities
A.2 Full Results